OpenStudy (anonymous):

A car travels 172 miles using 7 gallons of gas. At that rate, how far can the car travel using 42 gallons of gas A.1,032 miles B.1,548 miles C.1,204 miles D.1,376 miles

OpenStudy (anonymous):

\[\frac{ 172 }{ 7 } = \frac{ x }{ 42 }\]Set up the above ratio and solve for "x".

OpenStudy (anonymous):

It should be 1032 I think 172/7=24.57 24.57x42=1032

OpenStudy (texaschic101):

tcarroll is correct....set it up as a proportion, cross multiply, then solve for x

OpenStudy (anonymous):

thx guys now its my turn to solve it lol but if have any more method's than please tell me

OpenStudy (anonymous):

Staying away from decimals and simplifying, you get:\[\frac{ 172 \times 42 }{ 7 } = x = 172 \times 6\]

OpenStudy (anonymous):

thanks im not hot stuff in math but you guys are true friend's well like bro lol

OpenStudy (anonymous):

anyone good with graph's

OpenStudy (anonymous):

thx for the recognition, @Feesa . Good luck to you in all of your studies. I can handle graphs.